OpenStudy (anonymous):

Solve for x. |2x + 2| = 14 x = 6 and x = -6 x = 6 and x = -8 x = 8 and x = -8 x = -6 and x = 8

OpenStudy (anonymous):

6,-8

OpenStudy (anonymous):

solve two linear equations \[2x+2=14\] and \[2x+2=-14\] separately, you will get two different answers

OpenStudy (anonymous):

2x+2=14=6 -2x-2=14=-8

OpenStudy (anonymous):

not to be critical, but what does \(2x+2=14=6\) mean?

OpenStudy (anonymous):

okay 14-2=12/2=6

OpenStudy (anonymous):

oooh ok means \(2x+2=14\implies x=6\) got it!

OpenStudy (anonymous):

on the other hand... \(14-2\neq \frac{12}{2}\)